OpenStudy (riley):

I'm really stumped on this problem, need some help. =/ Use technology to find the quadratic regression curve through the given points. {(0, -1), (-3, 3), (-1, 1), (-5, 2)} y(x) = ? Would I use my graphing calculator for this? I'm trying to use Excel for it right now, but it's not working out very well. >_<

OpenStudy (amistre64):

is this a best fit curve?

OpenStudy (amistre64):

i havent had any experience in actually regressing data

OpenStudy (riley):

Ummm I think so, I'm not 100% sure what the problem is asking to be honest . . . which is why I'm so confused. >_<

OpenStudy (riley):

thats all it's asking is to find the quadratic regression curve through the points it has listed. I tried graphing it, but I have no idea what to do after that, and the graph looks very strange.

OpenStudy (riley):

The example in my book is doing the "best fits" curve . . . but not a very good job of explaining how -_-

OpenStudy (anonymous):

Using R > Ldat <- data.frame("X"=c(0,-3,-1,-5), "Y"=c(-1,3,1,2)) > (LM<-lm(formula = Y ~ X + I(X^2), data = Ldat)) Call: lm(formula = Y ~ X + I(X^2), data = Ldat) Coefficients: (Intercept) X I(X^2) -1.0201 -2.4209 -0.3631 Y = -0.3631 X² -2.4209 X -1.0201